OpenStudy (anonymous):

Solve: 0 = 6x2 − 10x − 4. x = −1 over 3, x = −2 x = −1 over 3, x = 2 x = 1 over 3, x = −2 x = 1 over 3, x = 2

OpenStudy (anonymous):

\[6x ^{2}-10x-4=0\] \[6x ^{2}-12x+2x-4=0\] \[6x(x-2)+2(x-2)=0\] \[(6x+2)(x-2)=0\] \[(6x+2)=0 \cup (x-2)=0\] \[x=\frac{ -1 }{ 3 } \cup x=2\]